OpenStudy (darkbluechocobo):

On a speed time graph, what is the acceleration during the interval if the velocity changes from 250 m/s to 50 m/s in 2 s

OpenStudy (darkbluechocobo):

-100 m/s^2 100m/s^2 250 m/s^2 50m/s^2

OpenStudy (darkbluechocobo):

its a 200 change s:

OpenStudy (darkbluechocobo):

in velocity at least

OpenStudy (kl0723):

in a graph it is calculated by the slope :P

OpenStudy (darkbluechocobo):

o.o how do you get the slope for this?

OpenStudy (kl0723):

they don't give you points but they give you the change... and acceleration if you can recall is the change of speed over time... so do that ;)

OpenStudy (darkbluechocobo):

so the time is 2 seconds and the change in speed is -150 soh -150/2?

OpenStudy (darkbluechocobo):

thats -75 T_T im wrong

OpenStudy (darkbluechocobo):

wait im an idiots thats why

OpenStudy (darkbluechocobo):

the change is 200

OpenStudy (darkbluechocobo):

-200/2

OpenStudy (darkbluechocobo):

so -100

OpenStudy (darkbluechocobo):

that is my final answer i believe

OpenStudy (darkbluechocobo):

so -100m/s^2 is the acceleration

OpenStudy (kl0723):

initial vel= 250 m/s final vel= 50 m/s times interval= 2 s change in velocity is (Vfinal - Vinitial)

OpenStudy (darkbluechocobo):

so (50-250)= -200

OpenStudy (kl0723):

you were right :P lol

OpenStudy (darkbluechocobo):

whoop i kinda was like wait 250-150 :I thats 100 and i facepalmed

OpenStudy (darkbluechocobo):

Aha do you mind going over more I dont wanna be like a bother aha

OpenStudy (kl0723):

go ahead :P

OpenStudy (darkbluechocobo):

hokay aha

OpenStudy (darkbluechocobo):

Only flat line segments on a speed time graph have zero acceleration

OpenStudy (darkbluechocobo):

is this true?

OpenStudy (kl0723):

well you can do this yourself by putting a bit of thinking into it...

OpenStudy (kl0723):

let me ask you... is speed changing? on a flat line?

OpenStudy (kl0723):



OpenStudy (darkbluechocobo):